# Deviations from random matrix entanglement statistics for kicked quantum chaotic spin-$1/2$ chains

### v1

@@ Abstract
It is commonly expected that for quantum chaotic many body systems the statistical properties approach those of random matrices when increasing the system size. We demonstrate for various kicked spin-$1/2$ chain models that the average eigenstate entanglement indeed approaches the random matrix result. However, the distribution of the eigenstate entanglement differs significantly. While for autonomous systems such deviations are expected, they are surprising for the more scrambling kicked systems. We attribute the origin of the deviations to the local two-dimensional Hilbert spaces. This is also supported by similar deviations occurring in a local random matrix model with global diagonal coupling.
@@
The characterization of statistical properties of complex quantum systems has become an important tool in many areas of physics. Such a statistical description has started with complex nuclei, which can be described by results from random matrix theory (RMT) GuhMueWei1998 ; Meh2004 . Even single-particle systems follow RMT if the corresponding classical dynamics is chaotic, as stated by the famous Bohigas-Giannoni-Schmit conjecture BohGiaSch1984 . Based on a semiclassical trace formula description Gut1990 , this can be explained for the spectral form factor using specific properties of periodic orbits Ber1985 ; SieRic2001 ; Sie2002 ; MueHeuBraHaaAlt2004 ; MueHeuBraHaaAlt2005 ; MueHeuAltBraHaa2009 . More recently, it has become common to call a many-body system quantum chaotic if its level spacing distribution follows the RMT prediction, see e.g. Refs. JenSha1985 ; HsuAng1993 ; MonPoiBelSir1993 ; JacShe1997 ; AviRicBer2002 ; San2004 ; PonPapHuvAba2015 ; KheLazMoeSon2016 ; AkiWalGutGuh2016 ; BorLueSchKnaBlo2017 ; LuiBar2017b ; KosLjuPro2018 . A general expectation is that for a quantum chaotic many body system, also other statistics, like eigenvector distributions or entanglement follow the RMT results Izr1990 ; ZelBroFraHor1996 ; AleKafPolRig2016 .
@@
For autonomous many-body systems states at the edge of the spectrum are well known to differ from RMT EisCraPle2010 and in particular show area law entanglement Sre1993 . However it has recently been found, that even the midspectrum eigenstates do not fulfil the RMT expectations. Deviations from the RMT results are found at the level of eigenvector distributions, such as the components of the eigenvectors SrdProSot2021 and the fractal dimensions BaeHaqKha2019 ; PauCarRodBuc2021 , but also for the entanglement entropy VidHacBiaRig2017 ; VidRig2017 ; BeuBaeMoeHaq2018 ; HaqClaKha2022 . Some of these deviations can be understood analytically Hua2019 ; Hua2021 . An intuitive understanding for the deviations is based on the orthogonality of the eigenstates which implies that the non-RMT states at the edge of the spectrum impose for local Hamiltonians an additional structure on the midspectrum eigenstates HaqClaKha2022 .
@@
The situation is much less understood for Floquet many-body systems such as kicked systems. These systems are more scrambling in the sense that the time evolution operator is less local than in case of autonomous systems. Their dynamics is described by an effective Hamiltonian which have interactions in all possible ranges and orders HerKieBae2023 . Additionally, there is no energy dependence of the statistical properties of eigenstates. Therefore, the theoretical understanding of the autonomous case does not carry over to the kicked situation. Recently it has been shown that for kicked quantum chaotic many-body systems spectral properties such as the spectral form factor are excellently described by RMT at least in the thermodynamic limit KosLjuPro2018 ; BerKosPro2018 ; ChaLucCha2018b ; ChaLucCha2018a ; FriChaDeCha2019 ; FlaBerPro2020 ; RoyPro2020 ; BerKosPro2021 ; GarCha2021 ; ChaDeCha2021 ; FriKie2024 . Also for some average statistical properties of eigenstates agreement with RMT predictions was shown AleRig2014 ; SieLewScaZak2023 .
@@
However, a detailed understanding of the entanglement behavior for quantum chaotic kicked spin chains is still lacking. In particular, will kicked spin chains approach the expected random matrix predictions in the thermodynamic limit of large dimension?
@@
In this paper we establish for the prototypical class of quantum chaotic kicked spin-chains that the average entanglement entropy and the standard deviation approach RMT predictions. However, surprisingly, the distribution of the von Neumann entropy does not approach the distribution for the corresponding RMT ensemble. The distributions stay close but distinct, even for large system sizes, which strongly suggests that this persists also in the thermodynamic limit. This effect also occurs for a simple random matrix model with global coupling. Thus the deviations are not caused by the short range interactions in the kicked spin chains, but by the two-dimensional local Hilbert spaces.
@@
Figure 2: Characterization of the distribution of the von Neumann entropy in depence on the chain length $L$ for different models. (a) Absolute difference between average values and the COE prediction, Eq. (8). (b) Absolute difference between standard deviations the numerically computed COE value, Eq. (9). (c) Ratio of the absolute difference between the average values and the standard deviation, Eq. (10). As models KFIM (black $+$), $C_{\text{NN}}$ chain (green crosses), $C_{\text{NNNN}}$ chain (blue circles), and random matrix model (yellow squares).
@@
Assuming two distributions with approaching average values and standard deviations three different situations can arise, as illustrated in Fig. 3. Figure 3(a) shows the case, for which the distributions actually approach each other. In (b) the distributions remain close but do not approach each other and in (c) the two distributions form distinct peaks. To distinguish between these three cases we use the ratio of the absolute distance between the average values to the standard deviation PauCarRodBuc2021
@@
We now apply this criterion to the KFIM. Figure. 2(c) shows the ratio (10) in dependence of the chain length $L$. Initially, the ratio (10) increases, while for large $L$ it appears to saturate with a value $\mathcal{R}\approx 1.25$ for $L=20$. This leads to the remarkable conclusion, that the distribution of the von Neumann entropy for the KFIM and the COE remain close but do not approach each other with increasing system size. This result is also supported by a similar behavior of the Kullback-Leibler-divergence KulLei1951 ; CovTho2006 and the Jensen-Shannon-divergence Lin1991 (not shown).
@@
and $H^{z}$ as for the KFIM, see Eq. (6). This model has an additional next–next–nearest neighbor coupling in the $x$ component. Additionally, we study a random matrix model for quantum chaotic Floquet many body systems given by FriKie2024 ,
@@
with $U_{j}$ for $j\in\{1,2,\dots,L\}$ being independent $2\times 2$ matrices chosen randomly from the CUE and $\xi_{k}$ for $k\in\{1,\dots,2^{L}\}$ being random variables uniformly distributed in $\left[-\pi,\pi\right[$. Similar to the spin-$1/2$ chain models, this model has two-dimensional local Hilbert spaces, but with a diagonal global coupling. Numerical paramters are given in Note1 .
@@
Since this behavior occurs for all models with very different coupling ranges, the origin for the deviations cannot be the local coupling. Instead, the cause must be a property common to all models. Therefore we conclude that the deviations are due to the local two-dimensional Hilbert spaces of the models. The type of coupling is responsible for differences in the details of the behavior of ratio (10) and its approach towards saturation.
Discussion and outlook – We demonstrate that for prototypical examples of quantum chaotic kicked spin-$1/2$ chains the average eigenstate entanglement approaches the COE result when increasing the system size. However, we find that the distribution of the entanglement entropy remains significantly different from the distribution for the COE. These deviations for the entropy are demonstrated for three different kicked spin-$1/2$ chain models and a random matrix model, which takes into account the local two-dimensional Hilbert spaces of the chains but contains a global coupling. Thus the observed deviations are due to the local substructure, while the coupling of the subsystems affect the form of the deviations.
We have also performed numerical computations of the linear entropy (closely related to the purity), for which similar deviations of the statistics from the results for the COE are found. Therefore one could think that also other eigenvector statistics might be affected. It turns out that the multifractal dimensions of eigenstates, see e.g. BaeHaqKha2019 ; PauCarBucRod2021 and references therein, for all considered models approach the COE results. This means, that entanglement statistics is more sensitive to deviations from RMT than the multifractal dimensions.
In the future it would be interesting to understand the influence of the dimensions of the subsystems. For example, for the random matrix model it is observed FriKie2024 that for large subsystems the spectral form factor approaches the CUE instead of the COE. Possible systems to study this would have a larger spin of the subsystems, e.g. spin-1 or spin-2 chains. Another important question is whether and if so to what extent the deviations of the eigenstate entanglement statistics from the COE influence the dynamical behavior of the system. Similar deviations of the entanglement of time-evolved wave-packets could mean that thermalization properties will be different from the RMT expectations.

### v2

@@ Abstract
It is commonly expected that for quantum chaotic many body systems the statistical properties approach those of random matrices when increasing the system size. We demonstrate for various kicked spin-$1/2$ chain models that the average eigenstate entanglement indeed approaches the random matrix result. However, the distribution of the eigenstate entanglement differs significantly. While for autonomous systems such deviations are expected, they are surprising for the more scrambling kicked systems. Similar deviations occur in a tensor-product random matrix model with all-to-all interactions. Therefore we attribute the origin of the deviations for the kicked spin-$1/2$ chain models to the tensor-product structure of the Hilbert spaces. As a consequence this would mean that such many body systems cannot be described by the standard random matrix ensembles.
@@
The characterization of statistical properties of complex quantum systems has become an important tool in many areas of physics. Such a statistical description has started with complex nuclei, which can be described by results from random matrix theory (RMT) GuhMueWei1998 ; Meh2004 . Even single-particle systems follow RMT if the corresponding classical dynamics is chaotic, as stated by the famous Bohigas-Giannoni-Schmit conjecture BohGiaSch1984 . Based on a semiclassical trace formula description Gut1990 , this can be explained for the spectral form factor using specific properties of periodic orbits Ber1985 ; SieRic2001 ; Sie2002 ; MueHeuBraHaaAlt2004 ; MueHeuBraHaaAlt2005 ; MueHeuAltBraHaa2009 . More recently, it has become common to call a many-body system quantum chaotic if its level spacing distribution follows the RMT prediction, see e.g. Refs. JenSha1985 ; HsuAng1993 ; MonPoiBelSir1993 ; JacShe1997 ; AviRicBer2002 ; San2004 ; PonPapHuvAba2015 ; KheLazMoeSon2016 ; AkiWalGutGuh2016 ; BorLueSchKnaBlo2017 ; LuiBar2017b ; KosLjuPro2018 . A general expectation is that for a quantum chaotic many-body systems, also other statistics, like eigenvector distributions or entanglement follow the RMT results Izr1990 ; ZelBroFraHor1996 ; AleKafPolRig2016 .
@@
For autonomous many-body systems states at the edge of the spectrum are well known to differ from RMT EisCraPle2010 and in particular show area law entanglement Sre1993 . However it has recently been found, that even the midspectrum eigenstates do not fulfil the RMT expectations. Deviations from the RMT results are found at the level of eigenvector distributions, such as the components of the eigenvectors SrdProSot2021 and the fractal dimensions BaeHaqKha2019 ; PauCarRodBuc2021 , but also for the entanglement entropy VidHacBiaRig2017 ; VidRig2017 ; BeuBaeMoeHaq2018 ; HaqClaKha2022 ; KliSwiVidRig2023 . Some of these deviations can be understood analytically Hua2019 ; Hua2021 . An intuitive understanding for the deviations is based on the orthogonality of the eigenstates which implies that the non-RMT states at the edge of the spectrum impose for local Hamiltonians an additional structure on the midspectrum eigenstates HaqClaKha2022 .
@@
The situation is much less understood for Floquet many-body systems such as kicked systems. These systems are more scrambling in the sense that the time evolution operator is less local than in case of autonomous systems. Their dynamics is described by an effective Hamiltonian which have interactions in all possible ranges and orders HerKieBae2023 . Additionally, there is no energy dependence of the statistical properties of eigenstates. Therefore, the theoretical understanding of the autonomous case does not carry over to the kicked situation. Recently it has been shown that for kicked quantum chaotic many-body systems spectral properties such as the spectral form factor are excellently described by RMT at least in the thermodynamic limit KosLjuPro2018 ; BerKosPro2018 ; ChaLucCha2018b ; ChaLucCha2018a ; FriChaDeCha2019 ; FlaBerPro2020 ; RoyPro2020 ; BerKosPro2021 ; GarCha2021 ; ChaDeCha2021 ; FriKie2024 . Also for some average statistical properties of eigenstates agreement with RMT predictions was shown AleRig2014 ; SieLewScaZak2023 . In contrast, it was recently found, that for the model of random Floquet circuits, the distribution of the entanglement entropy deviates from RMT even for large systems RodJonKhe2024 .
@@
However, a detailed understanding of the entanglement behavior for quantum chaotic kicked systems such as spin chains is still lacking. In particular, will kicked spin chains approach the expected random matrix predictions in the thermodynamic limit of large dimension?
@@
In this paper we establish for the prototypical class of quantum chaotic kicked spin-chains that the average entanglement entropy and the standard deviation approach RMT predictions. However, surprisingly, the distribution of the von Neumann entropy does not approach the distribution for the corresponding RMT ensemble. The distributions stay close but distinct, even for large system sizes, which strongly suggests that this persists also in the thermodynamic limit. This effect also occurs for a simple random matrix model with all-to-all interactions. This suggests that the deviations are not caused by the short range interactions in the kicked spin chains, but by the tensor-product structure of the Hilbert spaces.
@@
Figure 2: Characterization of the distribution of the von Neumann entropy in dependence on the chain length $L$ for different models. (a) Absolute difference between average values and the COE prediction, Eq. (8). (b) Absolute difference between standard deviations and the numerically computed COE value, Eq. (9). (c) Ratio of the absolute difference between the average values and the standard deviation, Eq. (10). Models: KFIM (black $+$), $C_{\text{NN}}$ chain (green crosses), $C_{\text{NNNN}}$ chain (blue circles), and random matrix model (yellow squares).
@@
Figure 4: Distribution of rescaled von Neumann entropy for the KFIM with different $L$ and a standard normal distribution (red dashed line) which provides a good approximation of the distribution for the COE.
@@
Assuming two distributions with approaching average values and standard deviations three different situations can arise, as illustrated in Fig. 3. Figure 3(a) shows the case, for which the distributions actually approach each other. In (b) the distributions remain close but do not approach each other and in (c) the two distributions form distinct peaks. To distinguish between these three cases we use the ratio of the absolute distance between the average values to the standard deviation PauCarRodBuc2021 ; RodJonKhe2024
@@
We now apply this criterion to the KFIM. Figure. 2(c) shows the ratio (10) in dependence of the chain length $L$. Initially, the ratio (10) increases, while for large $L$ it appears to saturate with a value $\mathcal{R}\approx 1.25$ for $L=20$. This leads to the remarkable conclusion, that the distribution of the von Neumann entropy for the KFIM and the COE remain close but do not approach each other with increasing system size. This result is also supported by a similar behavior of the Kullback-Leibler-divergence KulLei1951 ; CovTho2006 and the Jensen-Shannon-divergence Lin1991 (not shown). Another way to visualize the persistence of the deviations is to consider the rescaling $\chi=(S_{1}-\overline{S_{1}^{\text{COE}}})/\sigma(S_{1}^{\text{COE}})$. This is shown for the KFIM in comparison to the COE result in Fig. 4. One clearly sees, that there is no convergence towards COE.
@@
and $H^{z}$ as for the KFIM, see Eq. (6). This model has an additional next–next–nearest neighbor coupling in the $x$ component. Note that neither the $C_{\text{NN}}$ chain nor the $C_{\text{NNNN}}$ chain has a self-dual point. Additionally, we study a random matrix model for quantum chaotic Floquet many body systems given by FriKie2024 ,
@@
with $U_{j}$ for $j\in\{1,2,\dots,L\}$ being independent $2\times 2$ matrices chosen randomly from the CUE and $\xi_{k}$ for $k\in\{1,\dots,2^{L}\}$ being random variables uniformly distributed in $\left[-\pi,\pi\right[$. Similar to the spin-$1/2$ chain models, this model has a Hilbert space, which is the tensor-product of $L$ two-dimensional Hilbert spaces, but due to the all-to-all interaction lacks a spatial locality structure. Note that due to the coupling there is also no block diagonal structure in the resulting time evolution operator. Numerical parameters are given in Note1 .
@@
To confirm the generality of these results, we have also investigated spin-$1/2$ chains with next-nearest-neighbor coupling and long-range power-law coupling, as well as the case of unequal dimensions of the subsystems and chains with an odd number of spins (not shown). All these examples show the same effect. As a consequence we attribute the deviations of the entanglement statistics to the common feature of all these systems, which is the tensor-product structure of the Hilbert spaces. Note that the details of the behavior of ratio (10) and its approach towards saturation are system dependent.
Discussion and outlook – We demonstrate that for prototypical examples of quantum chaotic kicked spin-$1/2$ chains the average eigenstate entanglement approaches the COE result when increasing the system size. However, we find that the distribution of the entanglement entropy remains significantly different from the distribution for the COE. These deviations for the entropy are demonstrated for different kicked spin-$1/2$ chain models and a random matrix model, which takes into account the product structure of the Hilbert spaces of the chains but contains all-to-all interactions. This is in agreement with the results of Ref. RodJonKhe2024 for the model of random Floquet circuits which also show deviations from the full random matrix results and their origin is attributed to the locality of the model. We have also performed numerical computations of the linear entropy (closely related to the purity), for which similar deviations of the statistics from the results for the COE are found.
Combining these results we attribute the deviations to the underlying product structure of the Hilbert spaces. As a consequence this would mean that such many body systems cannot be described by the standard random matrix ensembles.
One could think that also other eigenvector statistics might be affected. It turns out that the multifractal dimensions of eigenstates, see e.g. BaeHaqKha2019 ; PauCarBucRod2021 and references therein, for all considered models follow the COE results. This means, that entanglement statistics is more sensitive to deviations from RMT than the multifractal dimensions.
In the future it would be interesting to understand the influence of the dimensions of the subsystems. Since we attribute the origin of the deviations to the product structure of the Hilbert spaces, we expect to find a similar effect also in chains with a larger spin of the subsystems, e.g. spin-1 or spin-$3/2$ chains. Another important question is whether and if so to what extent the deviations of the eigenstate entanglement statistics from the COE influence the dynamical behavior of the system. Similar deviations of the entanglement of time-evolved wave-packets could mean that thermalization properties will be different from the RMT expectations.